OpenStudy (anonymous):

The current in a solenoid with 23 turns per centimeter is 0.45A . The solenoid has a radius of 1.6cm . A long, straight wire runs along the axis of the solenoid, carrying a current of 15A . Part A Find the magnitude of the net magnetic field a radial distance of 0.80cm from the straight wire.

OpenStudy (anonymous):

@KingGeorge @TuringTest please help me with physics. I have tried all the questions but they come out wrong

OpenStudy (anonymous):

given, n (no. of turns per meter ) = 23 * 1000 = 23000 I' (current through the solenoid) = 0.45A R (radius of the solenoid) = 1.6 cm= 1.6/100 m I''(current through the straight wire) = 15A r (distance of the point from the wire) = 0.8 cm = 0.8/100 m Let B' be the magnetic field of the solenoid. \[B' = \mu nI' = 4 \pi * 10^{-7} * 23 * 10^{2} * 0.45\] B' = 1.29 * 10^-6 T Let B'' be the magnetic field of the long wire \[B'' = \frac{ \mu I'' }{ 2\pi r} = \frac{ 4\pi * 10^{-7} * 15 }{ 2\pi * 0.8 * 10 ^{-2} }\] B'' = 3.75 * 10^-4 T The total magnetic field is B = B' + B'' =1.29 * 10^-6 + 3.75 * 10^-4 = 3.7629 * 10^-4 T